please show work
4) (24 points) What is the solubility of Ag2SO4 (silver sulfate) in... a) ...pure water? b) ...1.5 M AgNO3?
Answer #1

Vuwer het sis molar Solublity of Ag2sou Solublity product of Ag, soy Ksp = 1.20x10-5 @ Pane walter Az soy as = 2 Ay + sou - 2low so s is since Ksp lu very also very small 28 << 1.5 1.20x10-5 = (1.5) (5) S: 5.33 X 10-6 molle Solublity = 3.33 Vlo-t mod